OpenStudy (anonymous):

calculate the number of milliliters of an aqueous .300M zinc nitrate solution required to provide 21.0 grams of zinc nitrate

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) Look out for the atomic weight in your periodic table for each element present in Zinc nitrate compound. Zn(NO3)2 Zn:65.4 N:14 O:16 21.0/65.4+(14*2)+(16*6)=21.0/189.4=0.110 moles Number of Moles=Concentration(mol/dm^3)*Volume 0.110/0.300=Volume Volume=0.369 dm^3 To change from dm^3 to milliliters we multiply by 1000 369 milliliters

OpenStudy (anonymous):

How did you get .369? I'm getting .366